2 Al + 3 CuCh2 -› 2 AICl, + 3 Cu


Molar masses: 27.0 g/mol Al,


134.5 g/mol CuCh2, 63.5 g/ mol


Cu


Question 1: How many grams of CuCh2 are


necessary to consume 10 g Al?


Question 2: How many grams of Cu will be


produced?

1
Expert's answer
2022-06-03T03:44:03-0400

According to the following reaction, when 2 mol (54 g) of Al reacts, 3 mol (405 g) of CuCl2 reacts to form 3 mol (192 g) of Cu.


2Al + 3CuCl2 = 2AlCl3 + 3Cu


1.

When 10 g of Al reacts, we find how many g of CuCl2 react:

54 g Al — 405 g CuCl2

10 g Al — X

X = 10 * 405 / 54 = 75 g CuCl2


2.

We find how many grams of Cu are formed when 10 g of Al reacts:

54 g Al — 192 g Cu

10 g Al — X

X = 10 * 192 / 54 = 35,56 g Cu